### v1

# Quantum Coherent Transport of 1D ballistic states in second order topological insulator Bi4Br4

Abstract
We investigate quantum transport in micrometer-sized single crystals of Bi4Br4, a material predicted to be a second-order topological insulator. 1D topological states with long phase coherence times are revealed via the modulation of quantum interferences with magnetic field and gate voltage. In particular, we demonstrate the existence of Aharanov-Bohm interference between 1D ballistic states several micrometers long, that we identify as phase-coherent hinge modes on neighbouring step edges at the crystal surface. These Aharanov-Bohm interferences are made possible by a disordered phase-coherent contact region, the existence of which is confirmed by STEM/EDX imaging of FIB lamellae. Their coherent nature modulates the transmission of the 1D edge states, leading to weak antilocalization and universal conductance fluctuations with surprisingly large characteristic fields and a strongly anisotropic behavior. These complementary experimental results provide a comprehensive, coherent description of quantum transport in Bi4Br4, and establishes the material as belonging to the class of second-order topological insulators with topologically protected 1D ballistic states.
I Introduction
An important recent achievement in condensed matter physics is the discovery of a new class of materials, Topological Insulators (TI), whose bulk is insulating, while the boundaries conduct current in a quasi-ideal way [1, 2, 3]. In particular, the 1D edges of 2DTIs realize the Quantum Spin Hall (QSH) state [4, 5], where current is carried dissipationlessly by two counter-propagating ballistic channels with opposite spins, forming a helical edge state. This opens many possibilities, ranging from dissipationless charge and spin transport at room temperature, to new avenues for quantum computing. Among the recent additions to the TI classification, Second Order Topological Insulators (SOTIs) [6, 7] are n-dimensional crystals exhibiting n-2 topological states. In particular, 3D SOTIs have insulating bulk and surfaces, but topologically protected, one-dimensional helical “hinge” states. 1D helical states have been evidenced in several systems in 2D such as HgTe quantum wells [8], monolayer WTe2 [9, 10], or TaIrTe4 [13] and in 3D systems such as crystalline bismuth [11, 12]. Unfortunately, the expected topological protection has turned out to be less robust than anticipated, notably due to the existence of conduction in the bulk and surface, or charge puddles [14]. This complicates the fundamental study of the edge states, and motivates the search for different TIs with a reduced contribution of the non-topological bulk states. Among newly discovered TIs, Bi4Br4 appears to be a very promising material, with a large bulk gap ( meV), and strong experimental indications of a SOTI character.
This material, which in its form has a bilayer structure of chains oriented along the crystal’s b direction (Fig. 1a,b), has indeed been predicted to show the Quantum Spin Hall state for the monolayer [15, 16] and has been classified as a SOTI in the multilayer limit [17, 18, 19, 21, 20]. The SOTI nature in the multilayer case results from the coupling of helical edge states upon stacking of 2D monolayers, leading to gapped side surfaces with a gap of the order of 30 meV. On a perfect crystal, only two helical states at hinges remain uncoupled, whose location depends on the parity of the number of layers.
The band structure of Bi4Br4 has been probed by ARPES [22, 23, 24] and micro-ARPES [22, 25]. This latter work confirmed the existence of gapped lateral surfaces, with a gap ranging from 28 to 40 meV, and resolved their spin texture, indicating a spin polarization perpendicular to the direction of the chains. Optical pump–probe microscopy at room temperature revealed a relatively long carrier lifetime for boundary states attributed to their linear dispersion relation [26]. Conducting hinge states have been probed by STM [27, 28, 23, 29] up to room temperature, on mono- and few-layers-high step edges. These states are gapped by a magnetic field perpendicular to chains with a Landé factor of 53, indicating a spin quantization axis along the b axis, in stark contrast to the spin polarization of bulk [20] and lateral surface states [25].
With regard to electronic transport, experiments carried out on macroscopic samples indicate that the temperature dependence of resistivity is close to that of a semiconductor below 50 K and often shows a local maximum around 200 K [31, 32, 33, 22, 30, 23, 34]. It has been proposed that the former results either from competition between bulk and hinge states [22, 30], or the appearance of a charge density wave [32, 33] or a variable-range hopping induced by Br vacancies [34]. The 200 K feature is attributed to a Lifshitz transition [30, 23]. At low temperature, macroscopic samples exhibit weak localization or weak antilocalization depending on the facet [35, 36]. High-pressure experiments found a metal-insulator transition that can even lead to a superconducting state at pressures above 3.8 GPa [33]. By probing micrometer-size crystals obtained by exfoliation, several temperature behaviors were obtained : a nearly temperature independent resistivity for relatively long flakes [37], a gate tunable resistance attributed to competition of bulk and surface conduction [38], or a thermal activation of bulk carriers [39]. This latter reference demonstrated an Aharonov-Bohm effect associated to quantum interference of coherent hinge states around single lateral surfaces in four- and six-layer flakes.
In the present work, we study quantum transport in micrometer-size Bi4Br4 flakes with thicknesses ranging from 50 to 150 nm. We find evidence of phase-coherent ballistic 1D states over several microns via Aharonov-Bohm effect, as well as negative 4-wire resistances. Interestingly, the samples also consistently display weak antilocalization and universal conductance fluctuations, which are staples of 2D coherent transport in the diffusive regime. These apparently conflicting observations can be reconciled by considering a small coherent disordered region between the pristine crystal and the contact, which enables both inter-hinge coherence and anomalous 2D quantum interference. The existence of this region is confirmed by direct imaging and elemental analysis using STEM/EDX. While this contact regime prevents observation of the quantification of conductance generally expected in QSH systems, it allows extremely rich signatures of the underlying topological transport.
II Exfoliation and contacting Bi4Br4 flakes
The Bi4Br4 crystals were synthesised by the chemical vapour transport method [40, 22]. The samples were exfoliated in a glove box under argon atmosphere with O0.2 ppm and H2O5 ppm. The exfoliation and transfer of m-size Bi4Br4 flakes was done using PMDS stamps. We check with scanning transmission electron microscope (STEM) imaging that the crystalline quality of the exfoliated flakes is preserved before any contacts is made [65].
In order to contact the flakes we developed two strategies. The first one is based on contacts made on top of the flake. This is done by e-beam lithography, with low temperature processes to preserve the quality of the flakes (PMMA resist baking below 100℃), and metal deposition in an e-beam evaporator after argon ion beam etching (IBE). Different contact materials were tried. The best results were obtained with palladium (Pd). The samples are thus contacted by Pd(7 nm)/Au(100 nm) electrodes, this leads to k range two-wire resistances at room temperature. However this method damages the crystal under the contacts and exposes the flakes to organic solvents and ambient air during the e-beam lithography and lift-off processes. To avoid these problems, we also tried contacting the flakes by transferring them onto predefined metallic electrodes, fabricated by e-beam and optical lithography, with e-beam metal deposition. The Bi4Br4 flakes are then pressed onto the electrodes in Ar atmosphere. During this process, we found that heating the transfer stage to 80°C produced ohmic contacts more reliably. The best results were obtained, as in the case of contacts on top, with palladium electrodes (7 nm thickness) and also yields two-wire resistances in the k range.
To understand the nature of the contact region between Pd electrodes and Bi4Br4 flakes, we produced FIB lamella of the region of the contacts, as well as slices away from the contact and analyzed them in a STEM. The results are presented in figure 1. Sections taken far from the contacts show excellent crystalline quality, even after exposure to ambient air (Fig. 1c,d). By contrast, for both contacting schemes, the contact regions are disordered, with Pd/Bi interdiffusion, and have a typical 100 nm size.
III Anisotropic electronic transport in flakes
We have performed electronic transport measurement on more than ten samples. In the following we focus on four samples. One, denoted , is a flake with six contacts (top contacts with Pd(7 nm)/Au(100 nm) electrodes), thus defining five segments, with length ranging from 1 m to 5 m. The flake has a total length of 30 m, a width of 2.1 m and a thickness of around 100 nm (Fig. 4a). The flake is deposited on a doped silicon substrate, which serves as a back gate, covered by 500 nm of oxide.
We will also discuss three samples with a Hall bar geometry. Sample , shown on figure 2a, is a flake deposited on top of Ti(5 nm)/Au(10 nm)/Pd(5 nm) contacts. Sample has similar contacts, while sample has contacts deposited from the top (Fig. 5 and [65]).
All samples display 4-wire resistance values in the k range at room temperature, that do not change significantly down to 10 K (Fig. 7a). There is a sharp increase of resistance below 1 K, which we attribute to the mesoscopic nature of our sample rather than to the band structure (discussed in the Supplementary information [65]). The absence of a strong temperature dependence in our measurements, down to 10 K, contrasts with other works [22, 30, 38, 32, 33, 34, 30, 23] and suggest that our samples are relatively clean, with a low density of doping defects, and therefore have a rather temperature-stable chemical potential. Moreover, the absence of a temperature-activated behavior across the whole temperature range suggests that the Fermi level lies at least 26 meV away from a bulk band edge, thus clearly in the bulk gap.
The 4-wire resistances of the segments sample do not increase proportionally to their lengths, which indicates that they are dominated by a contact resistance [65]. This can be explained by competing scenarios : either a resistive interface between the Pd/Au contacts and well conducting bulk or surface states, or ballistic conduction in the Landauer-Büttiker picture. The variations in resistance from segment to segment can thus be explained either by fluctuations in interface quality, or by differences in the number of edge modes contacted due to insulating resist residues under the electrodes.
In order to elucidate the flow of current inside the Hall bar devices, we use a formula derived in the context of graph theory [41, 42]:
which relates two quantities ; the Laplacian matrix and the resistance distance matrix , with a vector with all coefficients equal to 1. Modelling the sample as a network of lumped resistances with nodes representing terminals, the resistance distance matrix contains the 2-wire resistances between each possible pair of terminals of the resistance network. The Laplacian matrix has matrix elements , with the discrete node-to-node conductances, and diagonal coefficients ensuring current conservation . thus corresponds to the conductance matrix in the context of mesoscopic transport [43]. This result therefore provides a direct determination of the conductance matrix from simple resistance measurements.
The conductance matrix of is given on figure 2b (see [65] for samples and ). All contacts to the flake are doubled, so that the lead resistance can be removed for every 2-wire configuration. We find a very anisotropic transport, with relatively high conductances, of the order of the conductance quantum, along the direction of the crystal. Noticeably, the conduction along the direction is consistently among the smallest in all of our devices, with conductances less then 0.15 e²/h for contacts 200 nm apart for . Non-zero transverse conductances may result from residual interface resistances that are not eliminated by our measurement procedure.
Finally, we have carried out magnetotransport measurements (Fig. 3) on the same sample, which revealed an essentially symmetrical 4-wire transverse resistance dominated by quantum interferences at low temperature (see next section). The antisymmetrical fraction of the contribution to the magnetoresistance is extremely small, with no clear linear dependence in magnetic field, indicating the absence of Hall effect in the sample. This is consistent with the bulk and (001) surfaces of the Bi4Br4 crystal being insulating.
These results are a striking demonstration that the bulk and (001) surfaces are insulating. It suggests that conduction is due to lateral surfaces or is unidimensional along the b direction.
IV Low temperature quantum transport
At low temperature, all samples’ differential resistance exhibits a zero bias peak [65]. Considering the linear regime, with a low excitation current, and for temperatures below 20 K, magnetoconductance has two main characteristics: a weak-antilocalization (WAL) peak at zero magnetic field, and sizeable universal conductance fluctuations (UCF) (Fig. 4c for sample ).
The low-field WAL peaks are fitted using the Hikami-Larkin-Nagaoka relation :
with the digamma function, the magnetic field, the phase-coherent length and a prefactor corresponding to the number of independent channels [45, 44]. Equation 2 is valid for a 2D diffusive conductor in the limit of very strong-spin orbit coupling (spin-orbit length ). The fitting is done in the magnetic field range [-0.5T,0.5T] for T under 2 K, and [-3T,3T] otherwise with two adjustable parameters : the prefactor , which is kept constant for a given segment at all temperatures, and the phase-coherent length . Formula 2 gives an excellent agreement over the entire field range. This seems to indicate the absence of other contributions to the magnetoconductance. However fits using the WAL expression in the case of a quasi-1D conductors yielded both a poor agreement with the experimental data and inconsistent values for [65]. We find relatively short across all segments, between 100 nm and 200 nm at low temperatures (Fig. 4e) and values for between 2 and 8, which suggests the existence of several uncorrelated electron paths in parallel [44, 46, 47].
We now turn to the analysis of the UCFs, which were measured as a function of magnetic field (Fig. 4c) and gate voltage [65]. The UCFs exhibit a typical magnetic field scale, given by the half-width of the autocorrelation, which provides another determination of [48, 50, 49]. As shown in Fig. 4e, these values are close to those extracted from WAL, except at temperatures above 1K, for which the experimental noise outweighs the fluctuation signal.
For all sections of sample , , extracted from WAL and UCFs, follows a power-law dependence down to mK, below which it saturates. We find an exponent of , consistent with the expected for dephasing in 2D dominated by electron-electron interactions [44].
Despite being much shorter than the segment length, we find sizeable fluctuation amplitudes with no systematic decrease with increasing segment length. We recall that UCF in diffusive conductors should decrease as the segment length extends beyond , due to self-averaging. By contrast the absence of a systematic decrease in fluctuations amplitude with length (Fig. 4d) shows that the quantum interference we measure do not originate from diffusive path within the flake.
Rather, we propose that diffuse transport only occurs in the damaged region near the contacts, roughly 100 nm in size as discussed above. Those diffusive regions are connected by 1D ballistic channels, up to 5 m long, so that coherent transport occurs over up to 5 m. We conjecture that it is the transmission from contacts to ballistic channel that is modulated by quantum interference in the disordered regions. This is what leads to the absence of self-averaging and to the high field scale of UCF and WAL. This scenario was already pointed out in the context of the conductance of edge states in the integer quantum Hall effect [52] and is qualitatively reproduced by our numerical simulations (Figure 8a and [65]).
A confirmation of this picture is provided by the dependence of quantum interference on magnetic field orientation. We performed magnetoconductance measurements on samples using a vector magnet, allowing to apply magnetic fields in any directions up to 1T. We present field cuts in three different planes : in-plane, longitudinal and transverse, taken at 3.8 K to remove the contribution from conductance fluctuations, which complicates the analysis (Fig. 5a, b and c). The first two cuts clearly highlight a near absence of magnetoconductance along the axis of the flake (X direction of Fig. 5a and b). The longitudinal and transverse cut reveal that the decrease of conductance with magnetic field is maximum for out-of-plane field, leading an extended lobe in the polar plot in the perpendicular direction (indicated by the black dashed lines in Fig. 5), demonstrating that the surface of diffusion essentially lies in the plane of the sample. This behavior, together with the absence of self-averaging of the UCFs, rules out diffusion on the sides surfaces of Bi4Br4 as the source of the observed quantum interferences.
In addition to the large out-of-plane magnetoconductance observed in the transverse cut, several devices exhibit other lobes (gray dashed lines on Fig. 5c,f,g) that are not related to specific crystalline direction of Bi4Br4. These axes correspond to directions of the electrodes deposited on the lateral side of the flake. Indeed, the disordered region induced by the contacts follows the slope of the etched crystal, which defines the plane of diffusion for the WAL (Fig. 5e). Moreover, for sample , with a Hall bar geometry, we find that, for contacts on the side of the flakes, only one of the side lobes is present, and that it is more prominent than the out-of-plane one. This confirms that each side lobe comes from one flank of the flake. Samples contacted from below showed only an out-of-plane magnetoconductance, confirming that the origin of the side lobes lies in the contact method.
In conclusion the magneto-transport at low temperature and its anisotropy with the magnetic field orientation confirms that the regions determining the mesoscopic fluctuations are the diffusive contact regions, leading to a field dependent transmission to 1D ballistic channels.
V Negative 4-wire resistances
While the behavior of samples contacted from above and from below are similar on many aspects, there are a few clear differences. The first is the presence of a higher interface resistance for bottom contacts, and the second is the absence of a 4-wire negative differential resistance in top-contacted samples.
Indeed, we observe 4-wire negative resistance fluctuations that extend between -6 k and 10 k with a 2-wire resistance of 20-35 k of (Fig. 6). This corresponds to a fraction of the theoretical limit of , predicted for ballistic conduction measured with moderately invasive voltage probes [53]. Indeed, the measured 2-wire resistance of the voltage probe is high, indicating a weakly perturbative probes of the edge states that can favor the appearance of these negative fluctuations. Moreover, contacts from below may allow for the hinge states to be preserved to a certain extent, if they are not cut by a large Pd/Bi alloy region. We emphasize that this prediction of a 4-wire negative differential resistance for ballistic systems has not been reported by many studies. Experiments on single-walled carbon nanotubes have found negative fluctuations reaching the theoretical limit [54], as well as measurements on high-mobility GaAs 2DEGs, which were observed only in a specific field range [55, 56]. Both of these experiments featured exceptionally clean systems with non-invasive voltage probes, which contrasts with the disorder induced at the contacts with Bi4Br4. Therefore, this observation of negative differential resistance in our samples could be a consequence of the expected topological protection of hinge states to elastic disorder.
VI Aharonov-Bohm interferences
In addition to the aperiodic UCFs and WAL displayed in the magnetoconductance, several samples also display periodic oscillations, leading to well-defined peaks in their Fourier transform (FFT).
In particular the longitudinal magnetoresistance of sample shows a very clear periodicity of around 0.24 T at low temperatures leading to a FFT peak at 4 T-1 (Fig.7a and c). We interpret this as Aharonov-Bohm (AB) interferences [57] between two 1D states of length 6 m and separed by 2.7 nanometers, allowed by the partially coherent contact region, as confirmed by numerical simulations (Fig. 8b and [65]). One possible reason for this particularly clear signature of AB interference in this sample is the small separation between the lateral electrodes, which selects only a few channels that can carry current coherently across the flake without being interrupted by a contact (Fig.7b). We find an exponential decay with temperature of the AB amplitude (Inset of Fig. 7c). This temperature behaviour has been reported in quasi-1D ballistic rings [64], 3D topological insulators [60, 61, 62] and hinge state loops in thin MoTe2 flakes [63]. This dependence is however at odds with the result of Hossain et al. [39] on thin flakes of Bi4Br4, which instead exhibits a power law dependence that is attributed to the extremely long, in effect temperature-independent, phase coherence length of edge states. The functional difference between the decay of the UCFs and AB oscillations with temperature in our samples suggests that their coherence are limited by distinct mechanisms. One would expect the amplitude of AB oscillations to be proportional to the probability of an electron trajectory to complete a loop without phase-breaking event , with the perimeter of the closed trajectory [57]. The exponential decay of the peak amplitude in the FFT with temperature therefore implies a 1/T dependence of the phase coherence length and, taking m, m at 1 K. The fact that we can measure long phase coherence length implies that phase-breaking occurs predominantly within the hinge states, so that the electron trajectories contributing the most to the AB effect do not explore the diffusive regions much, thanks to the very small separation between hinge states (2.7 nm). However this result should be taken with caution since the AB oscillations are measured on the conductance of the combined system of two partially phase coherent disordered contacts (with 100 nm at low temperatures) and highly coherent edge states, whose interference effects are not entirely disentangled. Nevertheless our results emphasize the coherence of electronic transport in relatively long Bi4Br4 flakes.
Strikingly, for device , several of the FFT peaks correspond to a field scale proportional to the length of the considered segment. Hence if the magnetic frequency is rescaled by the segment length (Fig. 9b), it is possible to identify some peaks in the FFT of different segments that correspond to the same rescaled magnetic frequency. We interpret this as AB interferences around surfaces of fixed widths across each segment. We postulate that these interferences originate from coherent 1D hinge states that extend over the entire length of the flake. However they are divided into several segments by the fabricated contacts. Consequently, from segment to segment, there is a constant lateral separation between the hinge states, as illustrated in figure 9a. As the AB effect is visible for all segments, including the longest ones, the phase coherence length of the edge states must be of the order of 5 m or more. However, the limited phase coherence in the contacts region induces an upper cutoff to the lateral distance between 1D channels allowing for AB interferences, such that we do not observe peaks in the Fourier spectrum corresponding to distances greater than a few tens of nanometers. We attribute the absence of a clear magnetic field periodicity in the magneto-conductance signal to the multiplicity of frequencies present in the Fourier transform, which are due to the many 1D coherent channels of the flake that may interfere with each other. Moreover, the magnetic frequencies identified in the FFT spectrum can be related to some specific location of edge states, which are consistent with the crystalline structure of Bi4Br4 [65]. The temperature dependence of the identified peaks exhibits an exponential decay consistent with a 1/T dependence of phase coherence length and a value at 1K similar to the one of sample .
VII Conclusion
We have probed the electronic transport in micrometer-size Bi4Br4 flakes and have demonstrated a strongly anisotropic transport. We have highlighted two distinct types of contributions to low-temperature coherent transport and argue that they suggest of long coherent 1D hinge states contacted to reservoirs via a small coherent disordered region. The first type of contributions yields universal conductance fluctuations and weak-antilocalization and arises from phase coherence in the region close to the contacts, that modulates transmission through the 1D states. The second type of contributions is the periodic Aharonov-Bohm conductance oscillations due to interference between adjacent 1D coherent channels. We emphasize that these effects are due to the very nature of the contacts, which exhibit some phase coherence despite being rather invasive. Both effects indicate that 1D states are ballistic and coherent over several micrometers at low temperatures. The signature of the 1D ballistic states, located at step edges distributed at the surface of the Bi4Br4 crystal, and their interference are thus paradoxically revealed thanks to the coherent diffusive nature of the contact regions. These results argue in favor of crystal-preserving contact methods to achieve the conductance quantization regime in small samples with a reduced number of edge states. The signature of the helical nature of these 1D states in multilayer samples remains to be demonstrated. Furthermore, the use of superconducting contacts to enter the regime of proximity-induced superconductivity is of great interest, with expected perfect Andreev reflection, and would further enhance the potential of Bi4Br4 as a second-order topological insulator.
Acknowledgements.
The authors acknowledge B. Wieder, E. Gerber, T. Wakamura, M. Kociak, M. Aprili, J. Esteve for fruitful discussions, L. Bugaud’s involvement in the initial steps of the project and technical help from S. Autier-Laurent and R. Weil. This work was funded by JSPS KAKENHI (Grant Numbers 21H04652, 21K18181,and 21H05236), JST CREST (Grant Number JPMJCR20B4), the French Renatech network and the European Union through the BALLISTOP ERC 66566 advanced grant.References
- [1] M. Z. Hasan and C. L. Kane, Colloquium: Topological insulators, Rev. Mod. Phys. 82, 3045 (2010).
- [2] X.-L. Qi and S.-C. Zhang, Topological insulators and superconductors, Rev. Mod. Phys. 83, (2011).
- [3] Y. Ando, Topological Insulator Materials, J. Phys. Soc. Jpn. 82, 102001 (2013).
- [4] C. L. Kane and E. J. Mele, Topological Order and the Quantum Spin Hall Effect, Phys. Rev. Lett. 95, 146802 (2005).
- [5] B. A. Bernevig and S.-C. Zhang, Quantum Spin Hall Effect, Phys. Rev. Lett. 96, 106802 (2006).
- [6] F. Schindler, A. M. Cook, M. G. Vergniory, Z. Wang, S. S. P. Parkin, B. A. Bernevig, and T. Neupert, Higher-order topological insulators, Science Advances 4, eaat0346 (2018).
- [7] B. Xie, H.-X. Wang, X. Zhang, P. Zhan, J.-H. Jiang, M. Lu, and Y. Chen, Higher-order band topology, Nat Rev Phys 3, 520 (2021).
- [8] M. König, S. Wiedmann, C. Brüne, A. Roth, H. Buhmann, L. W. Molenkamp, X.-L. Qi, and S.-C. Zhang, Quantum Spin Hall Insulator State in HgTe Quantum Wells, Science 318, 766 (2007).
- [9] Z. Fei, T. Palomaki, S. Wu, W. Zhao, X. Cai, B. Sun, P. Nguyen, J. Finney, X. Xu, and D. H. Cobden, Edge conduction in monolayer WTe2, Nature Phys 13, 7 (2017).
- [10] S. Wu, V. Fatemi, Q. D. Gibson, K. Watanabe, T. Taniguchi, R. J. Cava, and P. Jarillo-Herrero, Observation of the quantum spin Hall effect up to 100 kelvin in a monolayer crystal, Science 359, 76 (2018).
- [11] A. Murani et al., Ballistic edge states in Bismuth nanowires revealed by SQUID interferometry, Nature Communications 8, ncomms15941 (2017).
- [12] F. Schindler et al., Higher-order topology in bismuth, Nature Phys 14, 918 (2018).
- [13] J. Tang et al., Dual quantum spin Hall insulator by density-tuned correlations in TaIrTe4, Nature 628, 515 (2024).
- [14] B. Weber et al., 2024 roadmap on 2D topological insulators, J. Phys. Mater. 7, 022501 (2024).
- [15] J.-J. Zhou, W. Feng, C.-C. Liu, S. Guan, and Y. Yao, Large-Gap Quantum Spin Hall Insulator in Single Layer Bismuth Monobromide Bi 4 Br 4, Nano Lett. 14, 4767 (2014).
- [16] J.-J. Zhou, W. Feng, G.-B. Liu, and Y. Yao, Topological edge states in single- and multi-layer Bi4Br4, New J. Phys. 17, 015004 (2015).
- [17] C.-H. Hsu, X. Zhou, Q. Ma, N. Gedik, A. Bansil, V. M. Pereira, H. Lin, L. Fu, S.-Y. Xu, and T.-R. Chang, Purely rotational symmetry-protected topological crystalline insulator -Bi4Br4, 2D Mater. 6, 031004 (2019).
- [18] F. Tang, H. C. Po, A. Vishwanath, and X. Wan, Efficient topological materials discovery using symmetry indicators, Nat. Phys. 15, 470 (2019).
- [19] C. Yoon, C.-C. Liu, H. Min, and F. Zhang, Quasi-One-Dimensional Higher-Order Topological Insulators, arXiv:2005.14710 [Cond-Mat] (2020).
- [20] K.-S. Lin et al., Spin-resolved topology and partial axion angles in three-dimensional insulators, Nat Commun 15, 550 (2024).
- [21] J. Han, W. Xiao, and Y. Yao, Quasi-one-dimensional topological material Bi4X4(X=Br,I), Advances in Physics: X 7, 2057234 (2022).
- [22] R. Noguchi et al., Evidence for a higher-order topological insulator in a three-dimensional material built from van der Waals stacking of bismuth-halide chains, Nature Materials 1 (2021).
- [23] Yang, M. et al. Large-Gap Quantum Spin Hall State and Temperature-Induced Lifshitz Transition in Bi 4 Br 4. ACS Nano 16, 3036–3044 (2022).
- [24] Zhong, J. et al. Coalescence of multiple topological orders in quasi-one-dimensional bismuth halide chains. Nat Commun 16, 1163 (2025).
- [25] W. Zhao et al., Topological electronic structure and spin texture of quasi-one-dimensional higher-order topological insulator Bi4Br4, Nat Commun 14, 8089 (2023).
- [26] J. Han et al., Optical bulk-boundary dichotomy in a quantum spin Hall insulator, Science Bulletin 68, 417 (2023).
- [27] N. Shumiya et al., Evidence of a room-temperature quantum spin Hall edge state in a higher-order topological insulator, Nat. Mater. 1 (2022).
- [28] J. K. Hofmann et al., Shear-Resistant Topology in Quasi One-Dimensional van Der Waals Material Bi4Br4, arXiv:2411.13320.
- [29] Peng, X. et al. Observation of Topological Edge States on Bi4Br4 Nanowires Grown on TiSe 2 Substrates. J. Phys. Chem. Lett. 12, 10465–10471 (2021).
- [30] D.-Y. Chen, D. Ma, J. Duan, D. Chen, H. Liu, J. Han, and Y. Yao, Quantum transport evidence of boundary states and Lifshitz transition in , Phys. Rev. B 106, 075206 (2022).
- [31] H. von Benda, A. Simon, and W. Bauhofer, Zur Kenntnis von BiBr und BiBr1,167, Zeitschrift für anorganische und allgemeine Chemie 438, 53 (1978).
- [32] T. G. Filatova, P. V. Gurin, L. Kloo, V. A. Kulbachinskii, A. N. Kuznetsov, V. G. Kytin, M. Lindsjo, and B. A. Popovkin, Electronic structure, galvanomagnetic and magnetic properties of the bismuth subhalides Bi4I4 and Bi4Br4, Journal of Solid State Chemistry 180, 1103 (2007).
- [33] X. Li et al., Pressure-induced phase transitions and superconductivity in a quasi–1-dimensional topological crystalline insulator -Bi4Br4, PNAS 116, 17696 (2019).
- [34] Z. Gong et al., Br-Vacancies Induced Variable Ranging Hopping Conduction in High-Order Topological Insulator Bi4Br4, Small Methods, 2400517 (2024).
- [35] J. Zhong, M. Yang, F. Ye, C. Liu, J. Wang, J. Wang, W. Hao, J. Zhuang, and Y. Du, Facet-dependent Electronic Quantum Diffusion in the High-Order Topological Insulator , Phys. Rev. Applied 17, 064017 (2022).
- [36] Zhong, J. et al. Observation of Anomalous Planar Hall Effect Induced by One-Dimensional Weak Antilocalization. ACS Nano acsnano.3c10120 (2024)
- [37] L. Qiao et al., Ultralong Single-Crystal -Bi4Br4 Nanobelts with a High Current Carrying Capacity by Mechanical Exfoliation, J. Phys. Chem. C 125, 22312 (2021).
- [38] S.-L. Wu, Z.-H. Ren, Y.-Q. Zhang, Y.-K. Li, J.-F. Han, J.-X. Duan, Z.-W. Wang, C.-Z. Li, and Y.-G. Yao, Gate-tunable transport in van der Waals topological insulator Bi4Br4 nanobelts, J. Phys.: Condens. Matter 35, 234001 (2023).
- [39] M. S. Hossain et al., Quantum transport response of topological hinge modes, Nat. Phys. 1 (2024).
- [40] R. Noguchi et al., A weak topological insulator state in quasi-one-dimensional bismuth iodide, Nature 566, 7745 (2019).
- [41] Bapat, R. B, Resistance matrix of a weighted graph, Communications in Mathematical and in Computer Chemistry 50, 73‑82 (2004).
- [42] Devriendt Karel, Graph Geometry from Effective Resistances, PhD Thesis, Trinity College, 2022.
- [43] Datta, S. Electronic Transport in Mesoscopic Systems. (Cambridge Univ. Press, Cambridge, 2009).
- [44] E. Akkermans and G. Montambaux, Mesoscopic Physics of Electrons and Photons, Cambridge University Press (2007).
- [45] S. Hikami, A. I. Larkin, and Y. Nagaoka, Spin-Orbit Interaction and Magnetoresistance in the Two Dimensional Random System, Progress of Theoretical Physics 63, 707 (1980).
- [46] H. Steinberg, J.-B. Laloë, V. Fatemi, J. S. Moodera, and P. Jarillo-Herrero, Electrically tunable surface-to-bulk coherent coupling in topological insulator thin films, Phys. Rev. B 84, 233101 (2011).
- [47] J. Chen, X. Y. He, K. H. Wu, Z. Q. Ji, L. Lu, J. R. Shi, J. H. Smet, and Y. Q. Li, Tunable surface conductivity in Bi 2 Se 3 revealed in diffusive electron transport, Phys. Rev. B 83, 241304 (2011).
- [48] P. A. Lee, A. D. Stone, and H. Fukuyama, Universal conductance fluctuations in metals: Effects of finite temperature, interactions, and magnetic field, Phys. Rev. B 35, 1039 (1987).
- [49] P. Mohanty and R. A. Webb, High-Field Measurements of Electron Decoherence Time in Metallic Nanowires: Switching off Magnetic Impurity Spins, Phys. Rev. Lett. 91, 066604 (2003).
- [50] C. W. J. Beenakker and H. Van Houten, Flux-cancellation effect on narrow-channel magnetoresistance fluctuations, Phys. Rev. B 37, 6544 (1988).
- [51] Groth, C. W., Wimmer, M., Akhmerov, A. R. and Waintal, X. Kwant: a software package for quantum transport. New J. Phys. 16, 063065 (2014).
- [52] M. Büttiker, Absence of backscattering in the quantum Hall effect in multiprobe conductors, Phys. Rev. B 38, 9375 (1988).
- [53] M. Büttiker, Symmetry of electrical conduction, IBM J. Res. Dev. 32, 317 (1988).
- [54] B. Gao, Y. F. Chen, M. S. Fuhrer, D. C. Glattli, and A. Bachtold, Four-Point Resistance of Individual Single-Wall Carbon Nanotubes, Phys. Rev. Lett. 95, 196802 (2005).
- [55] G. Timp, A. M. Chang, P. Mankiewich, R. Behringer, J. E. Cunningham, T. Y. Chang, and R. E. Howard, Quantum transport in an electron-wave guide, Phys. Rev. Lett. 59, 732 (1987).
- [56] de Picciotto, R., Stormer, H. L., Pfeiffer, L. N., Baldwin, K. W. and West, K. W. Four-terminal resistance of a ballistic quantum wire. Nature 411, 51–54 (2001).
- [57] S. Washburn and R. A. Webb, Aharonov-Bohm effect in normal metal quantum coherence and transport, Advances in Physics 35, 375 (1986).
- [58] Altshuler, B. L., Aronov, A. G. and Spivak, B. Z. The Aharonov–Bohm effect in disordered conductors. JETP Lett. 33, 94–97 (1981).
- [59] Washburn, S., Umbach, C. P., Laibowitz, R. B. and Webb, R. A. Temperature dependence of the normal-metal Aharonov–Bohm effect. Phys. Rev. B 32, 4789–4792 (1985).
- [60] Dufouleur, J. et al. Quasiballistic Transport of Dirac Fermions in a Bi 2 Se 3 Nanowire. Phys. Rev. Lett. 110, 186806 (2013).
- [61] Jauregui, L. A., Pettes, M. T., Rokhinson, L. P., Shi, L. and Chen, Y. P. Magnetic field-induced helical mode and topological transitions in a topological insulator nanoribbon. Nature Nanotech 11, 345–351 (2016).
- [62] G. Behner et al., Aharonov-Bohm Interference and Phase-Coherent Surface-State Transport in Topological Insulator Rings, Nano Lett. 23, 6347 (2023).
- [63] Pan, Z.-C. et al. Altshuler-Aronov-Spivak interference of one-dimensional helical edge states in . Phys. Rev. B 107, 045411 (2023).
- [64] A. E. Hansen, A. Kristensen, S. Pedersen, C. B. Sørensen, and P. E. Lindelof, Mesoscopic decoherence in Aharonov-Bohm rings, Phys. Rev. B 64, 045327 (2001).
- [65] See Supplemental Material for a discussion of the crystalline quality of Bi4Br4 flakes, resistance networks of two other samples at different temperatures, the length dependence of resistance for samples and , we expand on the temperature dependence of the Aharonov-Bohm effect for sample , details of the numerical simulations presented in the main text, the fitting of the magnetoconductance using the formula for 1D weak anti-localization, and finally a presentation of zero-bias anomaly data and comparison with theoretical predictions.

### v2

@@
-Quantum Coherent Transport of 1D ballistic states in second order topological insulator Bi4Br4
+Quantum Coherent Transport of 1D Ballistic States in Second-Order Topological Insulator Bi4Br4
 Abstract
 We investigate quantum transport in micrometer-sized single crystals of Bi4Br4, a material predicted to be a second-order topological insulator.
-1D topological states with long phase coherence times are revealed via the modulation of quantum interferences with magnetic field and gate voltage.
-In particular, we demonstrate the existence of Aharanov-Bohm interference between 1D ballistic states several micrometers long, that we identify as phase-coherent hinge modes on neighbouring step edges at the crystal surface.
-These Aharanov-Bohm interferences are made possible by a disordered phase-coherent contact region, the existence of which is confirmed by STEM/EDX imaging of FIB lamellae.
+1D topological states with long phase coherence times are revealed via the modulation of quantum interference with magnetic field and gate voltage.
+In particular, we demonstrate the existence of Aharonov-Bohm interference between 1D ballistic states several micrometers long, that we identify as phase-coherent hinge modes on neighboring step edges at the crystal surface.
+These Aharonov-Bohm oscillations are made possible by a disordered phase-coherent contact region, the existence of which is confirmed by scanning transmission electron microscopy combined with energy-dispersive X-ray spectroscopy (STEM-EDX) of FIB lamellae.
 Their coherent nature modulates the transmission of the 1D edge states, leading to weak antilocalization and universal conductance fluctuations with surprisingly large characteristic fields and a strongly anisotropic behavior.
-These complementary experimental results provide a comprehensive, coherent description of quantum transport in Bi4Br4, and establishes the material as belonging to the class of second-order topological insulators with topologically protected 1D ballistic states.
+These complementary experimental results provide a comprehensive, coherent description of quantum transport in Bi4Br4, and establish the material as a second-order topological insulator with topologically protected 1D ballistic states.
 I Introduction
-An important recent achievement in condensed matter physics is the discovery of a new class of materials, Topological Insulators (TI), whose bulk is insulating, while the boundaries conduct current in a quasi-ideal way [1, 2, 3].
-In particular, the 1D edges of 2DTIs realize the Quantum Spin Hall (QSH) state [4, 5], where current is carried dissipationlessly by two counter-propagating ballistic channels with opposite spins, forming a helical edge state.
+An important recent achievement in condensed matter physics is the discovery of a new class of materials, topological insulators (TIs), whose bulk is insulating, while the boundaries are conducting [1, 2, 3].
+In particular, the 1D edges of 2DTIs realize the quantum spin hall (QSH) state [4, 5], where current is carried dissipationlessly by two counterpropagating ballistic channels with opposite spins, forming a helical edge state.
 This opens many possibilities, ranging from dissipationless charge and spin transport at room temperature, to new avenues for quantum computing.
-Among the recent additions to the TI classification, Second Order Topological Insulators (SOTIs) [6, 7] are n-dimensional crystals exhibiting n-2 topological states.
+Among the recent additions to the TI classification, second order topological insulators (SOTIs) [6, 7] are -dimensional crystals exhibiting -dimensional topological states.
 In particular, 3D SOTIs have insulating bulk and surfaces, but topologically protected, one-dimensional helical “hinge” states.
@@
 Unfortunately, the expected topological protection has turned out to be less robust than anticipated, notably due to the existence of conduction in the bulk and surface, or charge puddles [14].
-This complicates the fundamental study of the edge states, and motivates the search for different TIs with a reduced contribution of the non-topological bulk states.
-Among newly discovered TIs, Bi4Br4 appears to be a very promising material, with a large bulk gap ( meV), and strong experimental indications of a SOTI character.
-This material, which in its form has a bilayer structure of chains oriented along the crystal’s b direction (Fig.
-1a,b), has indeed been predicted to show the Quantum Spin Hall state for the monolayer [15, 16] and has been classified as a SOTI in the multilayer limit [17, 18, 19, 21, 20].
-The SOTI nature in the multilayer case results from the coupling of helical edge states upon stacking of 2D monolayers, leading to gapped side surfaces with a gap of the order of 30 meV.
-On a perfect crystal, only two helical states at hinges remain uncoupled, whose location depends on the parity of the number of layers.
-The band structure of Bi4Br4 has been probed by ARPES [22, 23, 24] and micro-ARPES [22, 25].
+This complicates the fundamental study of the edge states, and motivates the search for different TIs with a reduced contribution of the nontopological bulk states.
+Among newly discovered TIs, Bi4Br4 appears to be a very promising material, with a large bulk gap ( meV) [15, 16, 17, 18, 19, 21, 20], and strong experimental indications of a SOTI character [22, 23, 24, 25] .
+This material, which in its form has a bilayer structure of chains oriented along the crystal’s direction [Fig.
+1(a) and 1(b)], has indeed been predicted to show the QSH state for the monolayer [15, 16] and the multilayer has been classified as a SOTI [17, 18, 19, 21, 20, 26, 27, 28, 23, 29].
+The SOTI character in the multilayer case results from the coupling of helical edge states upon stacking of 2D monolayers, leading to gapped side surfaces with a gap of the order of 30 meV.
+On a perfect crystal, only two helical states located at hinges remain uncoupled, which location depends on the parity of the number of layers.
+The band structure of Bi4Br4 has been probed by angle-resolved photoemission spectroscopy (ARPES) [22, 23, 24] and micro-ARPES [25].
 This latter work confirmed the existence of gapped lateral surfaces, with a gap ranging from 28 to 40 meV, and resolved their spin texture, indicating a spin polarization perpendicular to the direction of the chains.
-Optical pump–probe microscopy at room temperature revealed a relatively long carrier lifetime for boundary states attributed to their linear dispersion relation [26].
-Conducting hinge states have been probed by STM [27, 28, 23, 29] up to room temperature, on mono- and few-layers-high step edges.
-These states are gapped by a magnetic field perpendicular to chains with a Landé factor of 53, indicating a spin quantization axis along the b axis, in stark contrast to the spin polarization of bulk [20] and lateral surface states [25].
+In addition, optical pump–probe microscopy at room temperature revealed a relatively long carrier lifetime for boundary states attributed to their linear dispersion relation [26].
+Lastly, conducting hinge states have been probed by STM [27, 28, 23, 29] up to room temperature, on mono- and few-layers-high step edges.
+These states are gapped by a magnetic field perpendicular to chains with a Landé factor of 53, indicating a spin quantization axis along the axis, in stark contrast to the spin polarization of bulk [20] and lateral surface states [25].
 With regard to electronic transport, experiments carried out on macroscopic samples indicate that the temperature dependence of resistivity is close to that of a semiconductor below 50 K and often shows a local maximum around 200 K [31, 32, 33, 22, 30, 23, 34].
-It has been proposed that the former results either from competition between bulk and hinge states [22, 30], or the appearance of a charge density wave [32, 33] or a variable-range hopping induced by Br vacancies [34].
-The 200 K feature is attributed to a Lifshitz transition [30, 23].
-At low temperature, macroscopic samples exhibit weak localization or weak antilocalization depending on the facet [35, 36].
+It has been proposed that the former results from either competition between bulk and hinge states [22, 30], or the appearance of a charge density wave [32, 33] or a variable-range hopping induced by Br vacancies [34], while the 200 K feature is attributed to a Lifshitz transition [30, 23].
+At low temperature, macroscopic samples exhibit weak localization or weak antilocalization depending on the field orientation [35, 36].
 High-pressure experiments found a metal-insulator transition that can even lead to a superconducting state at pressures above 3.8 GPa [33].
-By probing micrometer-size crystals obtained by exfoliation, several temperature behaviors were obtained : a nearly temperature independent resistivity for relatively long flakes [37], a gate tunable resistance attributed to competition of bulk and surface conduction [38], or a thermal activation of bulk carriers [39].
+By probing micrometer-size crystals obtained by exfoliation, several temperature behaviors were obtained : a nearly temperature-independent resistivity for relatively long flakes [37], a gate tunable resistance attributed to competition of bulk and surface conduction [38], or a thermal activation of bulk carriers [39].
 This latter reference demonstrated an Aharonov-Bohm effect associated to quantum interference of coherent hinge states around single lateral surfaces in four- and six-layer flakes.
-In the present work, we study quantum transport in micrometer-size Bi4Br4 flakes with thicknesses ranging from 50 to 150 nm.
-We find evidence of phase-coherent ballistic 1D states over several microns via Aharonov-Bohm effect, as well as negative 4-wire resistances.
+In the present work, we study quantum transport in micrometer-sized Bi4Br4 flakes with thicknesses ranging from 50 to 150 nm.
+We find evidence of phase-coherent ballistic 1D states over several microns via the Aharonov-Bohm effect, as well as negative four-wire resistances.
 Interestingly, the samples also consistently display weak antilocalization and universal conductance fluctuations, which are staples of 2D coherent transport in the diffusive regime.
-These apparently conflicting observations can be reconciled by considering a small coherent disordered region between the pristine crystal and the contact, which enables both inter-hinge coherence and anomalous 2D quantum interference.
-The existence of this region is confirmed by direct imaging and elemental analysis using STEM/EDX.
-While this contact regime prevents observation of the quantification of conductance generally expected in QSH systems, it allows extremely rich signatures of the underlying topological transport.
-II Exfoliation and contacting Bi4Br4 flakes
-The Bi4Br4 crystals were synthesised by the chemical vapour transport method [40, 22].
+These apparently conflicting observations can be reconciled by considering a small, coherent, but disordered region between the pristine crystal and the contact, which enables both interhinge coherence and anomalous 2D quantum interference.
+The existence of this region is confirmed by direct imaging and elemental analysis using STEM-EDX.
+While this contact regime prevents observation of the quantification of conductance generally expected in QSH systems, it provides extremely rich signatures of the underlying topological transport.
+II Exfoliation and contacting of Bi4Br4 flakes
+The Bi4Br4 crystals were synthesized by the chemical vapor transport method [40, 22].
 The samples were exfoliated in a glove box under argon atmosphere with O0.2 ppm and H2O5 ppm.
-The exfoliation and transfer of m-size Bi4Br4 flakes was done using PMDS stamps.
-We check with scanning transmission electron microscope (STEM) imaging that the crystalline quality of the exfoliated flakes is preserved before any contacts is made [65].
+The exfoliation and transfer of micrometer-size Bi4Br4 flakes was done using PDMS stamps.
+We check with STEM imaging that the crystalline quality of the exfoliated flakes is preserved before any contacts is made (Sec.
+1 in Supplemental Material [41]).
 In order to contact the flakes we developed two strategies.
 The first one is based on contacts made on top of the flake.
-This is done by e-beam lithography, with low temperature processes to preserve the quality of the flakes (PMMA resist baking below 100℃), and metal deposition in an e-beam evaporator after argon ion beam etching (IBE).
+This is done by e-beam lithography, with low-temperature processes to preserve the quality of the flakes (PMMA resist baking below 100℃), and metal deposition in an e-beam evaporator after argon ion beam etching.
 Different contact materials were tried.
 The best results were obtained with palladium (Pd).
-The samples are thus contacted by Pd(7 nm)/Au(100 nm) electrodes, this leads to k range two-wire resistances at room temperature.
-However this method damages the crystal under the contacts and exposes the flakes to organic solvents and ambient air during the e-beam lithography and lift-off processes.
+The samples are thus contacted by Pd(7 nm)/Au(100 nm) electrodes; this leads to k two-wire resistances at room temperature.
+However, this method damages the crystal under the contacts and exposes the flakes to organic solvents and ambient air during the e-beam lithography and lift-off processes.
 To avoid these problems, we also tried contacting the flakes by transferring them onto predefined metallic electrodes, fabricated by e-beam and optical lithography, with e-beam metal deposition.
@@
 During this process, we found that heating the transfer stage to 80°C produced Ohmic contacts more reliably.
-The best results were obtained, as in the case of contacts on top, with palladium electrodes (7 nm thickness) and also yields two-wire resistances in the k range.
-To understand the nature of the contact region between Pd electrodes and Bi4Br4 flakes, we produced FIB lamella of the region of the contacts, as well as slices away from the contact and analyzed them in a STEM.
-The results are presented in figure 1.
-Sections taken far from the contacts show excellent crystalline quality, even after exposure to ambient air (Fig.
-1c,d).
-By contrast, for both contacting schemes, the contact regions are disordered, with Pd/Bi interdiffusion, and have a typical 100 nm size.
-III Anisotropic electronic transport in flakes
-We have performed electronic transport measurement on more than ten samples.
-In the following we focus on four samples.
-One, denoted , is a flake with six contacts (top contacts with Pd(7 nm)/Au(100 nm) electrodes), thus defining five segments, with length ranging from 1 m to 5 m.
-The flake has a total length of 30 m, a width of 2.1 m and a thickness of around 100 nm (Fig.
-4a).
+The best results were obtained, as in the case of contacts on top, with palladium electrodes and also yields two-wire resistances in the k range.
+To understand the nature of the contact region between Pd electrodes and Bi4Br4 flakes, we produced FIB lamellae of the region of the contacts, as well as slices away from the contacts and analyzed them in a STEM.
+The results are presented in Fig.
+1.
+Sections taken far from the contacts show excellent crystalline quality, even after exposure to ambient air [Fig.
+1(c) and 1(d)].
+By contrast, for both contacting schemes, the contact regions are disordered, with Pd/Bi interdiffusion, and have a typical size of approximately 100 nm.
+III Anisotropic electronic transport in Bi4Br4 flakes
+We have performed electronic transport measurements on more than ten samples.
+In the following, we focus on four samples.
+One, denoted , is a flake with six contacts [top contacts with Pd(7 nm)/Au(100 nm) electrodes], thus defining five segments, with length ranging from 1 to 5 m.
+The flake has a total length of 30 m, a width of 2.1 m and a thickness of around 100 nm [Fig.
+4(a)].
 The flake is deposited on a doped silicon substrate, which serves as a back gate, covered by 500 nm of oxide.
-We will also discuss three samples with a Hall bar geometry.
-Sample , shown on figure 2a, is a flake deposited on top of Ti(5 nm)/Au(10 nm)/Pd(5 nm) contacts.
-Sample has similar contacts, while sample has contacts deposited from the top (Fig.
-5 and [65]).
-All samples display 4-wire resistance values in the k range at room temperature, that do not change significantly down to 10 K (Fig.
-7a).
-There is a sharp increase of resistance below 1 K, which we attribute to the mesoscopic nature of our sample rather than to the band structure (discussed in the Supplementary information [65]).
-The absence of a strong temperature dependence in our measurements, down to 10 K, contrasts with other works [22, 30, 38, 32, 33, 34, 30, 23] and suggest that our samples are relatively clean, with a low density of doping defects, and therefore have a rather temperature-stable chemical potential.
+We also discuss three samples with a Hall bar geometry.
+Sample , shown in Fig.
+2(a), is a flake deposited on top of Ti(5 nm)/Au(10 nm)/Pd(5 nm) contacts.
+Sample has similar contacts, while sample has Pd(7 nm)/Au(100 nm) contacts deposited from the top (Fig.
+5 and Sec.
+2 of Supplemental Material [41]).
+All samples display four-wire resistance values in the k range at room temperature that do not change significantly down to 10 K [Fig.
+3(a)].
+There is a sharp increase of resistance below 1 K, which we attribute to the mesoscopic nature of our sample rather than to the band structure (Sec.
+7 of Supplemental Material [41]).
+The absence of a strong temperature dependence in our measurements contrasts with other works [22, 30, 38, 32, 33, 34, 30, 23] and suggests that our samples are relatively clean, with a low density of doping defects, and, therefore, have a rather temperature-stable chemical potential.
 Moreover, the absence of a temperature-activated behavior across the whole temperature range suggests that the Fermi level lies at least 26 meV away from a bulk band edge, thus clearly in the bulk gap.
-The 4-wire resistances of the segments sample do not increase proportionally to their lengths, which indicates that they are dominated by a contact resistance [65].
-This can be explained by competing scenarios : either a resistive interface between the Pd/Au contacts and well conducting bulk or surface states, or ballistic conduction in the Landauer-Büttiker picture.
-The variations in resistance from segment to segment can thus be explained either by fluctuations in interface quality, or by differences in the number of edge modes contacted due to insulating resist residues under the electrodes.
+The four-wire resistances of the segments of sample do not increase proportionally to their lengths, which indicates that they are dominated by a contact resistance (Sec.
+3 of Supplemental Material [41]).
+This can be explained by two competing scenarios : either a resistive interface between the Pd/Au contacts and well-conducting bulk or surface states, or ballistic conduction, likely due to hinge states.
+The variations in resistance from segment to segment can, thus, be explained either by fluctuations in interface quality or by differences in the number of edge modes contacted due to insulating resist residues under the electrodes.
 In order to elucidate the flow of current inside the Hall bar devices, we use a formula derived in the context of graph theory [63, 64]:
-which relates two quantities ; the Laplacian matrix and the resistance distance matrix , with a vector with all coefficients equal to 1.
-Modelling the sample as a network of lumped resistances with nodes representing terminals, the resistance distance matrix contains the 2-wire resistances between each possible pair of terminals of the resistance network.
+which relates two quantities, the Laplacian matrix and the resistance distance matrix , with a vector with all coefficients equal to 1.
+Modeling the sample as a network of lumped resistances with nodes representing terminals, the resistance distance matrix contains the two-wire resistances between each possible pair of terminals of the resistance network.
 The Laplacian matrix has matrix elements , with the discrete node-to-node conductances, and diagonal coefficients ensuring current conservation .
-thus corresponds to the conductance matrix in the context of mesoscopic transport [43].
-This result therefore provides a direct determination of the conductance matrix from simple resistance measurements.
-The conductance matrix of is given on figure 2b (see [65] for samples and ).
-All contacts to the flake are doubled, so that the lead resistance can be removed for every 2-wire configuration.
+, thus, corresponds to the conductance matrix in the context of mesoscopic transport [65].
+This result, therefore, provides a direct determination of the conductance matrix from simple resistance measurements [66].
+The conductance matrix of is given in Fig.
+2(b) (see Sec.
+2 in Supplemental Material [41] for samples and ).
+All contacts to the flake are doubled, so that the lead resistance can be removed for every two-wire configuration.
 We find a very anisotropic transport, with relatively high conductances, of the order of the conductance quantum, along the direction of the crystal.
-Noticeably, the conduction along the direction is consistently among the smallest in all of our devices, with conductances less then 0.15 e²/h for contacts 200 nm apart for .
-Non-zero transverse conductances may result from residual interface resistances that are not eliminated by our measurement procedure.
+Noticeably, the conduction along the direction is consistently among the smallest in all of our devices, with conductances under 0.15 for contacts 500 nm apart for .
+Nonzero transverse conductances may result from residual interface resistances that are not eliminated by our measurement procedure.
 Finally, we have carried out magnetotransport measurements (Fig.
-3) on the same sample, which revealed an essentially symmetrical 4-wire transverse resistance dominated by quantum interferences at low temperature (see next section).
+3) on the same sample, which revealed an essentially symmetrical four-wire transverse resistance dominated by quantum interferences at low temperature (see the next section).
 The antisymmetrical fraction of the contribution to the magnetoresistance is extremely small, with no clear linear dependence in magnetic field, indicating the absence of Hall effect in the sample.
-This is consistent with the bulk and (001) surfaces of the Bi4Br4 crystal being insulating.
-These results are a striking demonstration that the bulk and (001) surfaces are insulating. It suggests that conduction is due to lateral surfaces or is unidimensional along the b direction.
-IV Low temperature quantum transport
-At low temperature, all samples’ differential resistance exhibits a zero bias peak [65].
-Considering the linear regime, with a low excitation current, and for temperatures below 20 K, magnetoconductance has two main characteristics: a weak-antilocalization (WAL) peak at zero magnetic field, and sizeable universal conductance fluctuations (UCF) (Fig.
-4c for sample ).
+Accordingly, the longitudinal resistance exhibits no Shubnikov-de Haas oscillations (Sec.
+8 in Supplemental Material [41]), further demonstrating the absence of transverse transport.
+These results are a striking demonstration that the bulk and (001) surfaces are insulating. It suggests that conduction is either due to the lateral (100) surfaces or to 1D states along the direction.
+IV Low-temperature quantum transport
+At low temperature, the differential resistance of all samples exhibits a zero-bias peak, which cannot be attributed to self-heating (Sec.
+7 of Supplemental Material [41]).
+Considering the linear regime, with a low excitation current, and for temperatures below 20 K, magnetoconductance has two main characteristics: a weak-antilocalization (WAL) peak at zero magnetic field, and sizable universal conductance fluctuations (UCFs) [Fig.
+4(c) for sample ].
 The low-field WAL peaks are fitted using the Hikami-Larkin-Nagaoka relation :
-with the digamma function, the magnetic field, the phase-coherent length and a prefactor corresponding to the number of independent channels [45, 44].
-Equation 2 is valid for a 2D diffusive conductor in the limit of very strong-spin orbit coupling (spin-orbit length ).
-The fitting is done in the magnetic field range [-0.5T,0.5T] for T under 2 K, and [-3T,3T] otherwise with two adjustable parameters : the prefactor , which is kept constant for a given segment at all temperatures, and the phase-coherent length .
-Formula 2 gives an excellent agreement over the entire field range.
+with the digamma function, the magnetic field, the phase coherence length and a prefactor corresponding to the number of independent channels in parallel [68, 67].
+Equation (2) is valid for a 2D diffusive conductor in the limit of very strong spin-orbit coupling (spin-orbit length ).
+The fitting is done in the magnetic field range [-1 T, 1 T] for temperatures under 2 K, and [-3 T, 3 T] otherwise with two adjustable parameters : the prefactor , which is kept constant for a given segment at all temperatures, and the phase coherence length .
+Formula (2) gives an excellent agreement over the entire field range.
 This seems to indicate the absence of other contributions to the magnetoconductance.
-However fits using the WAL expression in the case of a quasi-1D conductors yielded both a poor agreement with the experimental data and inconsistent values for [65].
-We find relatively short across all segments, between 100 nm and 200 nm at low temperatures (Fig.
-4e) and values for between 2 and 8, which suggests the existence of several uncorrelated electron paths in parallel [44, 46, 47].
-We now turn to the analysis of the UCFs, which were measured as a function of magnetic field (Fig.
-4c) and gate voltage [65].
+However, fits using the WAL expression in the case of quasi-1D conductors yielded both a poor agreement with the experimental data and inconsistent values for (Sec.
+6 of Supplemental Material [41]).
+We find relatively short across all segments, between 100 nm and 200 nm at low temperatures [Fig.
+4(e)] and values for between 2 and 8, which suggests the existence of several uncorrelated electron paths in parallel [67, 69, 70].
+We now turn to the analysis of the UCFs, which were measured as a function of magnetic field [Fig.
+4(c)] and gate voltage (Sec.
+3 of Supplemental Material [41]).
 The UCFs exhibit a typical magnetic field scale, given by the half-width of the autocorrelation, which provides another determination of [71, 73, 72].
 As shown in Fig.
-4e, these values are close to those extracted from WAL, except at temperatures above 1K, for which the experimental noise outweighs the fluctuation signal.
-For all sections of sample , , extracted from WAL and UCFs, follows a power-law dependence down to mK, below which it saturates.
+4(e), these values are close to those extracted from WAL, except at temperatures above 1 K, for which the experimental noise outweighs the UCF signal.
+For all sections of sample , extracted from both WAL and UCFs follows a power-law dependence down to approximately 100 mK, below which it saturates.
 We find an exponent of , consistent with the expected for dephasing in 2D dominated by electron-electron interactions [67].
-Despite being much shorter than the segment length, we find sizeable fluctuation amplitudes with no systematic decrease with increasing segment length.
+Despite being much shorter than the segment length, we find sizable fluctuation amplitudes with no systematic decrease with increasing segment length.
 We recall that UCF in diffusive conductors should decrease as the segment length extends beyond , due to self-averaging.
-By contrast the absence of a systematic decrease in fluctuations amplitude with length (Fig.
-4d) shows that the quantum interference we measure do not originate from diffusive path within the flake.
-Rather, we propose that diffuse transport only occurs in the damaged region near the contacts, roughly 100 nm in size as discussed above.
+By contrast, the absence of a systematic decrease in fluctuations amplitude with length [Fig.
+4(d)] shows that the quantum interference we measure does not originate from a diffusive path within the flake.
+Rather, we propose that diffuse transport occurs only in the damaged region near the contacts, roughly 100 nm in size as discussed above.
 Those diffusive regions are connected by 1D ballistic channels, up to 5 m long, so that coherent transport occurs over up to 5 m.
@@
 This is what leads to the absence of self-averaging and to the high field scale of UCF and WAL.
-This scenario was already pointed out in the context of the conductance of edge states in the integer quantum Hall effect [52] and is qualitatively reproduced by our numerical simulations (Figure 8a and [65]).
-A confirmation of this picture is provided by the dependence of quantum interference on magnetic field orientation.
-We performed magnetoconductance measurements on samples using a vector magnet, allowing to apply magnetic fields in any directions up to 1T.
-We present field cuts in three different planes : in-plane, longitudinal and transverse, taken at 3.8 K to remove the contribution from conductance fluctuations, which complicates the analysis (Fig.
-5a, b and c).
-The first two cuts clearly highlight a near absence of magnetoconductance along the axis of the flake (X direction of Fig.
-5a and b).
-The longitudinal and transverse cut reveal that the decrease of conductance with magnetic field is maximum for out-of-plane field, leading an extended lobe in the polar plot in the perpendicular direction (indicated by the black dashed lines in Fig.
+This scenario was already pointed out in the context of the conductance of edge states in the integer quantum Hall effect [75] and is qualitatively reproduced by our numerical simulations [Fig.
+8(a) and section 5 in Supplemental Material [41]).
+To confirm this picture, we studied the dependence of quantum interference on magnetic field orientation.
+We performed magnetoconductance measurements on samples using a vector magnet, allowing it to apply magnetic fields in arbitrary directions up to 1 T.
+We present field cuts in three different planes : in-plane, longitudinal and transverse, all taken at 3.8 K to remove the contribution from conductance fluctuations, which complicates the analysis [Fig.
+5(a)-5(c)].
+The first two cuts clearly highlight a near absence of magnetoconductance along the axis of the flake [ direction of Figs.
+5(a) and 5(b)].
+The longitudinal and transverse cuts reveal that the decrease of conductance with magnetic field is maximum for an out-of-plane field, leading an extended lobe in the polar plot in the perpendicular direction (indicated by the red lines in Fig.
 5), demonstrating that the surface of diffusion essentially lies in the plane of the sample.
 This behavior, together with the absence of self-averaging of the UCFs, rules out diffusion on the sides surfaces of Bi4Br4 as the source of the observed quantum interferences.
-In addition to the large out-of-plane magnetoconductance observed in the transverse cut, several devices exhibit other lobes (gray dashed lines on Fig.
-5c,f,g) that are not related to specific crystalline direction of Bi4Br4.
-These axes correspond to directions of the electrodes deposited on the lateral side of the flake.
-Indeed, the disordered region induced by the contacts follows the slope of the etched crystal, which defines the plane of diffusion for the WAL (Fig.
-5e).
-Moreover, for sample , with a Hall bar geometry, we find that, for contacts on the side of the flakes, only one of the side lobes is present, and that it is more prominent than the out-of-plane one.
+In addition to the large out-of-plane magnetoconductance observed in the transverse cut, several devices contacted from the top exhibit other lobes [gray dashed lines in Figs.
+5(c), 5(f) and 5(g)] that are not related to any specific crystalline direction of Bi4Br4.
+Rather, these axes correspond to the planes of the electrodes deposited on the lateral side of the flake.
+Indeed, the disordered region induced by the contacts follows the slope of the etched crystal, which defines the plane of diffusion for the WAL [Fig.
+5(e)].
+Moreover, for sample , with a Hall bar geometry, we find that, for contacts on the side of the flakes, only one of the side lobes is present and that it is more prominent than the out-of-plane one.
 This confirms that each side lobe comes from one flank of the flake.
-Samples contacted from below showed only an out-of-plane magnetoconductance, confirming that the origin of the side lobes lies in the contact method.
-In conclusion the magneto-transport at low temperature and its anisotropy with the magnetic field orientation confirms that the regions determining the mesoscopic fluctuations are the diffusive contact regions, leading to a field dependent transmission to 1D ballistic channels.
-V Negative 4-wire resistances
+In contrast, samples contacted from below showed only an out-of-plane magnetoconductance, confirming that the origin of the side lobes lies in the contact method.
+In conclusion, magnetotransport at low temperature and its anisotropy with magnetic field orientation confirms that the regions determining the mesoscopic fluctuations are the diffusive contact regions, leading to a field-dependent transmission to 1D ballistic channels.
+V Negative four-wire resistance
 While the behavior of samples contacted from above and from below are similar on many aspects, there are a few clear differences.
-The first is the presence of a higher interface resistance for bottom contacts, and the second is the absence of a 4-wire negative differential resistance in top-contacted samples.
-Indeed, we observe 4-wire negative resistance fluctuations that extend between -6 k and 10 k with a 2-wire resistance of 20-35 k of (Fig.
+The first is the presence of a higher interface resistance for bottom contacts, and the second is the absence of a four-wire negative differential resistance in top-contacted samples.
+Indeed, we observe negative four-wire resistance fluctuations that extend between -6 and 10 k with a two-wire resistance of 20-35 k for (Fig.
 6).
-This corresponds to a fraction of the theoretical limit of , predicted for ballistic conduction measured with moderately invasive voltage probes [53].
-Indeed, the measured 2-wire resistance of the voltage probe is high, indicating a weakly perturbative probes of the edge states that can favor the appearance of these negative fluctuations.
-Moreover, contacts from below may allow for the hinge states to be preserved to a certain extent, if they are not cut by a large Pd/Bi alloy region.
-We emphasize that this prediction of a 4-wire negative differential resistance for ballistic systems has not been reported by many studies.
+This corresponds to a sizable fraction of the theoretical limit of , predicted for ballistic conduction measured with moderately invasive voltage probes [76].
+Indeed, the two-wire resistance measured through the voltage probes indicated in blue in Fig.
+6 is high (approximately 150 k), which may indicate a weakly perturbative contact to the edge states that can favor the appearance of these negative fluctuations.
+Indeed, contacts from below may allow for the hinge states to be preserved to a certain extent, if they are not cut by a large Pd/Bi alloy region (Fig.
+1).
+We emphasize that this prediction of a four-wire negative differential resistance for ballistic systems has not been reported by many studies.
 Experiments on single-walled carbon nanotubes have found negative fluctuations reaching the theoretical limit [77], as well as measurements on high-mobility GaAs 2DEGs, which were observed only in a specific field range [78, 79].
-Both of these experiments featured exceptionally clean systems with non-invasive voltage probes, which contrasts with the disorder induced at the contacts with Bi4Br4.
-Therefore, this observation of negative differential resistance in our samples could be a consequence of the expected topological protection of hinge states to elastic disorder.
-VI Aharonov-Bohm interferences
+Both of these experiments featured exceptionally clean systems with noninvasive voltage probes, which contrasts with the disorder induced at the contacts with Bi4Br4.
+Therefore, this observation of negative differential resistance in our samples could be a consequence of the expected topological protection of hinge states against elastic disorder.
+VI Aharonov-Bohm interference
 In addition to the aperiodic UCFs and WAL displayed in the magnetoconductance, several samples also display periodic oscillations, leading to well-defined peaks in their Fourier transform (FFT).
-In particular the longitudinal magnetoresistance of sample shows a very clear periodicity of around 0.24 T at low temperatures leading to a FFT peak at 4 T-1 (Fig.7a and c).
-We interpret this as Aharonov-Bohm (AB) interferences [57] between two 1D states of length 6 m and separed by 2.7 nanometers, allowed by the partially coherent contact region, as confirmed by numerical simulations (Fig.
-8b and [65]).
-One possible reason for this particularly clear signature of AB interference in this sample is the small separation between the lateral electrodes, which selects only a few channels that can carry current coherently across the flake without being interrupted by a contact (Fig.7b).
-We find an exponential decay with temperature of the AB amplitude (Inset of Fig.
-7c).
-This temperature behaviour has been reported in quasi-1D ballistic rings [64], 3D topological insulators [60, 61, 62] and hinge state loops in thin MoTe2 flakes [63].
-This dependence is however at odds with the result of Hossain et al.
-[39] on thin flakes of Bi4Br4, which instead exhibits a power law dependence that is attributed to the extremely long, in effect temperature-independent, phase coherence length of edge states.
-The functional difference between the decay of the UCFs and AB oscillations with temperature in our samples suggests that their coherence are limited by distinct mechanisms.
+In particular the longitudinal magnetoresistance of sample shows a very clear periodicity of around 0.24 T at low temperatures leading to a FFT peak at 4 T-1 [Fig.7(a) and 7(c)].
+Note that the periodicity of the oscillations does not depend on temperature but the amplitude does (Sec.
+4 of Supplemental Material [41]).
+We interpret this as Aharonov-Bohm (AB) interference [80] between two 1D states of length 6 m and separated by 2.7 nm, allowed by the partially coherent contact region, as confirmed by numerical simulations [Fig.
+8(b) and Sec.
+5 of Supplemental Material [41]).
+One possible reason for this particularly clear signature of AB interference in this sample is the small separation between the lateral electrodes, which selects only a few channels that can carry current coherently across the flake without being interrupted by a contact [Fig.7(b)].
+The good visibility of the AB oscillations is also favored by the small lateral distance between edge states, comparable to the mean free path in the diffusive contact [Sec.
+5 of Supplemental Material [41]].
+Besides, we find an exponential decay with temperature of the AB amplitude [inset in Fig.
+7(c)).
+This temperature behavior has been reported in quasi-1D ballistic rings [85], 3D topological insulators [81, 82, 83] and hinge state loops in thin MoTe2 flakes [84].
+This dependence is, however, at odds with the result of Hossain et al.
+[39] on thin flakes of Bi4Br4, which instead exhibits a power-law dependence that is attributed to the extremely long, effectively temperature-independent, phase coherence length of edge states.
+The functional difference between the decay of the UCFs and AB oscillations with temperature in our samples suggests that their coherence is limited by distinct mechanisms.
 One would expect the amplitude of AB oscillations to be proportional to the probability of an electron trajectory to complete a loop without phase-breaking event , with the perimeter of the closed trajectory [80].
-The exponential decay of the peak amplitude in the FFT with temperature therefore implies a 1/T dependence of the phase coherence length and, taking m, m at 1 K.
-The fact that we can measure long phase coherence length implies that phase-breaking occurs predominantly within the hinge states, so that the electron trajectories contributing the most to the AB effect do not explore the diffusive regions much, thanks to the very small separation between hinge states (2.7 nm).
-However this result should be taken with caution since the AB oscillations are measured on the conductance of the combined system of two partially phase coherent disordered contacts (with 100 nm at low temperatures) and highly coherent edge states, whose interference effects are not entirely disentangled.
-Nevertheless our results emphasize the coherence of electronic transport in relatively long Bi4Br4 flakes.
+The exponential decay of the peak amplitude in the FFT with temperature, therefore, implies a dependence of the phase coherence length and taking m, m at 1 K.
+The fact that we can measure such a long phase coherence length implies that phase breaking occurs predominantly within the hinge states, so that the electron trajectories contributing the most to the AB effect do not explore the diffusive regions much, thanks to the very small separation between hinge states (2.7 nm).
+Our results emphasize the coherence of electronic transport in relatively long Bi4Br4 flakes.
 Strikingly, for device , several of the FFT peaks correspond to a field scale proportional to the length of the considered segment.
-Hence if the magnetic frequency is rescaled by the segment length (Fig.
-9b), it is possible to identify some peaks in the FFT of different segments that correspond to the same rescaled magnetic frequency.
-We interpret this as AB interferences around surfaces of fixed widths across each segment.
-We postulate that these interferences originate from coherent 1D hinge states that extend over the entire length of the flake.
-However they are divided into several segments by the fabricated contacts.
-Consequently, from segment to segment, there is a constant lateral separation between the hinge states, as illustrated in figure 9a.
+Hence, if the magnetic frequency is rescaled by the segment length [Fig.
+9(b)], it is possible to identify some peaks in the FFT of different segments that correspond to the same rescaled magnetic frequency.
+We interpret this as AB interference around surfaces of fixed widths across each segment.
+We postulate that these oscillations originate from coherent 1D hinge states that extend over the entire length of the flake, which were subsequently divided into several segments by the fabricated contacts.
+Consequently, from segment to segment, there is a constant lateral separation between the hinge states, as illustrated in Fig.
+9(a).
+In fact, the magnetic frequencies identified in the FFT spectrum can be related to some specific configurations of hinge states that are consistent with the crystalline structure of Bi4Br4 [Fig.
+9(b)].
 As the AB effect is visible for all segments, including the longest ones, the phase coherence length of the edge states must be of the order of 5 m or more.
-However, the limited phase coherence in the contacts region induces an upper cutoff to the lateral distance between 1D channels allowing for AB interferences, such that we do not observe peaks in the Fourier spectrum corresponding to distances greater than a few tens of nanometers.
-We attribute the absence of a clear magnetic field periodicity in the magneto-conductance signal to the multiplicity of frequencies present in the Fourier transform, which are due to the many 1D coherent channels of the flake that may interfere with each other.
-Moreover, the magnetic frequencies identified in the FFT spectrum can be related to some specific location of edge states, which are consistent with the crystalline structure of Bi4Br4 [65].
-The temperature dependence of the identified peaks exhibits an exponential decay consistent with a 1/T dependence of phase coherence length and a value at 1K similar to the one of sample .
+However, the limited phase coherence in the contact regions induces an upper cutoff to the lateral distance between 1D channels allowing for AB interferences, such that we do not observe peaks in the Fourier spectrum corresponding to distances greater than a few tens of nanometers.
+Moreover, we attribute the absence of a clear magnetic field periodicity in the magnetoconductance signal to the multiplicity of frequencies present in the Fourier transform, which are due to the many 1D coherent channels of the flake that may interfere with each other.
+The temperature dependence of the identified peaks exhibits an exponential decay consistent with a dependence of phase coherence length and a value at 1 K similar to the one of sample .
+Moreover the existence of AB interference in all segments of up to 5 µm is consistent with long phase coherence length.
 VII Conclusion
 We have probed the electronic transport in micrometer-size Bi4Br4 flakes and have demonstrated a strongly anisotropic transport.
-We have highlighted two distinct types of contributions to low-temperature coherent transport and argue that they suggest of long coherent 1D hinge states contacted to reservoirs via a small coherent disordered region.
-The first type of contributions yields universal conductance fluctuations and weak-antilocalization and arises from phase coherence in the region close to the contacts, that modulates transmission through the 1D states.
-The second type of contributions is the periodic Aharonov-Bohm conductance oscillations due to interference between adjacent 1D coherent channels.
-We emphasize that these effects are due to the very nature of the contacts, which exhibit some phase coherence despite being rather invasive.
-Both effects indicate that 1D states are ballistic and coherent over several micrometers at low temperatures.
+We have highlighted two distinct types of contributions to low-temperature coherent transport and argue that they reveal the existence of long coherent 1D hinge states contacted to reservoirs via a small coherent disordered region.
+The first type of contribution yields universal conductance fluctuations and weak antilocalization, arising from phase coherence in the region close to the contacts, which modulates transmission through the 1D states.
+The second type of contribution is the periodic Aharonov-Bohm conductance oscillations due to interference between adjacent 1D coherent channels.
+We emphasize that both of these effects are due to the very nature of the contacts, which exhibit some phase coherence despite being rather invasive.
+These effects indicate that the 1D states are ballistic and coherent over several micrometers at low temperatures.
 The signature of the 1D ballistic states, located at step edges distributed at the surface of the Bi4Br4 crystal, and their interference are thus paradoxically revealed thanks to the coherent diffusive nature of the contact regions.
-These results argue in favor of crystal-preserving contact methods to achieve the conductance quantization regime in small samples with a reduced number of edge states.
-The signature of the helical nature of these 1D states in multilayer samples remains to be demonstrated.
-Furthermore, the use of superconducting contacts to enter the regime of proximity-induced superconductivity is of great interest, with expected perfect Andreev reflection, and would further enhance the potential of Bi4Br4 as a second-order topological insulator.
+It would be interesting to achieve the opposite limit of clean, crystal-preserving contacts in small samples with a reduced number of edge states, in order to access the conductance quantization regime.
+Also, the signature of the helical nature of these 1D states in multilayer samples remains to be demonstrated.
+Lastly, the use of superconducting contacts to enter the regime of proximity-induced superconductivity is of great interest, with expected perfect Andreev reflection, and would further enhance the potential of Bi4Br4 as a second-order topological insulator.
 Acknowledgements.
@@
 Kociak, M.
-Aprili, J.
+Aprili and J.
 Esteve for fruitful discussions, L.
-Bugaud’s involvement in the initial steps of the project and technical help from S.
+Bugaud’s involvement in the initial steps of the project, and technical help from S.
 Autier-Laurent and R.
 Weil.
-This work was funded by JSPS KAKENHI (Grant Numbers 21H04652, 21K18181,and 21H05236), JST CREST (Grant Number JPMJCR20B4), the French Renatech network and the European Union through the BALLISTOP ERC 66566 advanced grant.References
-- [1] M. Z. Hasan and C. L. Kane, Colloquium: Topological insulators, Rev. Mod. Phys. 82, 3045 (2010).
-- [2] X.-L. Qi and S.-C. Zhang, Topological insulators and superconductors, Rev. Mod. Phys. 83, (2011).
-- [3] Y. Ando, Topological Insulator Materials, J. Phys. Soc. Jpn. 82, 102001 (2013).
-- [4] C. L. Kane and E. J. Mele, Topological Order and the Quantum Spin Hall Effect, Phys. Rev. Lett. 95, 146802 (2005).
-- [5] B. A. Bernevig and S.-C. Zhang, Quantum Spin Hall Effect, Phys. Rev. Lett. 96, 106802 (2006).
-- [6] F. Schindler, A. M. Cook, M. G. Vergniory, Z. Wang, S. S. P. Parkin, B. A. Bernevig, and T. Neupert, Higher-order topological insulators, Science Advances 4, eaat0346 (2018).
-- [7] B. Xie, H.-X. Wang, X. Zhang, P. Zhan, J.-H. Jiang, M. Lu, and Y. Chen, Higher-order band topology, Nat Rev Phys 3, 520 (2021).
-- [8] M. König, S. Wiedmann, C. Brüne, A. Roth, H. Buhmann, L. W. Molenkamp, X.-L. Qi, and S.-C. Zhang, Quantum Spin Hall Insulator State in HgTe Quantum Wells, Science 318, 766 (2007).
-- [9] Z. Fei, T. Palomaki, S. Wu, W. Zhao, X. Cai, B. Sun, P. Nguyen, J. Finney, X. Xu, and D. H. Cobden, Edge conduction in monolayer WTe2, Nature Phys 13, 7 (2017).
-- [10] S.
-Wu, V.
-Fatemi, Q.
-D.
-Gibson, K.
-Watanabe, T.
-Taniguchi, R.
-J.
-Cava, and P.
-Jarillo-Herrero, Observation of the quantum spin Hall effect up to 100 kelvin in a monolayer crystal, Science 359, 76 (2018).
-- [11] A. Murani et al., Ballistic edge states in Bismuth nanowires revealed by SQUID interferometry, Nature Communications 8, ncomms15941 (2017).
-- [12] F. Schindler et al., Higher-order topology in bismuth, Nature Phys 14, 918 (2018).
-- [13] J. Tang et al., Dual quantum spin Hall insulator by density-tuned correlations in TaIrTe4, Nature 628, 515 (2024).
-- [14] B. Weber et al., 2024 roadmap on 2D topological insulators, J. Phys. Mater. 7, 022501 (2024).
-- [15] J.-J. Zhou, W. Feng, C.-C. Liu, S. Guan, and Y. Yao, Large-Gap Quantum Spin Hall Insulator in Single Layer Bismuth Monobromide Bi 4 Br 4, Nano Lett. 14, 4767 (2014).
-- [16] J.-J. Zhou, W. Feng, G.-B. Liu, and Y. Yao, Topological edge states in single- and multi-layer Bi4Br4, New J. Phys. 17, 015004 (2015).
-- [17] C.-H.
-Hsu, X.
-Zhou, Q.
-Ma, N.
-Gedik, A.
-Bansil, V.
-M.
-Pereira, H.
-Lin, L.
-Fu, S.-Y.
-Xu, and T.-R.
-Chang, Purely rotational symmetry-protected topological crystalline insulator -Bi4Br4, 2D Mater.
-6, 031004 (2019).
-- [18] F. Tang, H. C. Po, A. Vishwanath, and X. Wan, Efficient topological materials discovery using symmetry indicators, Nat. Phys. 15, 470 (2019).
-- [19] C. Yoon, C.-C. Liu, H. Min, and F. Zhang, Quasi-One-Dimensional Higher-Order Topological Insulators, arXiv:2005.14710 [Cond-Mat] (2020).
-- [20] K.-S. Lin et al., Spin-resolved topology and partial axion angles in three-dimensional insulators, Nat Commun 15, 550 (2024).
-- [21] J. Han, W. Xiao, and Y. Yao, Quasi-one-dimensional topological material Bi4X4(X=Br,I), Advances in Physics: X 7, 2057234 (2022).
-- [22] R. Noguchi et al., Evidence for a higher-order topological insulator in a three-dimensional material built from van der Waals stacking of bismuth-halide chains, Nature Materials 1 (2021).
-- [23] Yang, M. et al. Large-Gap Quantum Spin Hall State and Temperature-Induced Lifshitz Transition in Bi 4 Br 4. ACS Nano 16, 3036–3044 (2022).
-- [24] Zhong, J. et al. Coalescence of multiple topological orders in quasi-one-dimensional bismuth halide chains. Nat Commun 16, 1163 (2025).
-- [25] W. Zhao et al., Topological electronic structure and spin texture of quasi-one-dimensional higher-order topological insulator Bi4Br4, Nat Commun 14, 8089 (2023).
-- [26] J. Han et al., Optical bulk-boundary dichotomy in a quantum spin Hall insulator, Science Bulletin 68, 417 (2023).
-- [27] N. Shumiya et al., Evidence of a room-temperature quantum spin Hall edge state in a higher-order topological insulator, Nat. Mater. 1 (2022).
-- [28] J. K. Hofmann et al., Shear-Resistant Topology in Quasi One-Dimensional van Der Waals Material Bi4Br4, arXiv:2411.13320.
-- [29] Peng, X. et al. Observation of Topological Edge States on Bi4Br4 Nanowires Grown on TiSe 2 Substrates. J. Phys. Chem. Lett. 12, 10465–10471 (2021).
-- [30] D.-Y. Chen, D. Ma, J. Duan, D. Chen, H. Liu, J. Han, and Y. Yao, Quantum transport evidence of boundary states and Lifshitz transition in , Phys. Rev. B 106, 075206 (2022).
-- [31] H. von Benda, A. Simon, and W. Bauhofer, Zur Kenntnis von BiBr und BiBr1,167, Zeitschrift für anorganische und allgemeine Chemie 438, 53 (1978).
-- [32] T.
-G.
-Filatova, P.
-V.
-Gurin, L.
-Kloo, V.
-A.
-Kulbachinskii, A.
-N.
-Kuznetsov, V.
-G.
-Kytin, M.
-Lindsjo, and B.
-A.
-Popovkin, Electronic structure, galvanomagnetic and magnetic properties of the bismuth subhalides Bi4I4 and Bi4Br4, Journal of Solid State Chemistry 180, 1103 (2007).
-- [33] X. Li et al., Pressure-induced phase transitions and superconductivity in a quasi–1-dimensional topological crystalline insulator -Bi4Br4, PNAS 116, 17696 (2019).
-- [34] Z. Gong et al., Br-Vacancies Induced Variable Ranging Hopping Conduction in High-Order Topological Insulator Bi4Br4, Small Methods, 2400517 (2024).
-- [35] J.
-Zhong, M.
-Yang, F.
-Ye, C.
-Liu, J.
-Wang, J.
-Wang, W.
-Hao, J.
-Zhuang, and Y.
-Du, Facet-dependent Electronic Quantum Diffusion in the High-Order Topological Insulator , Phys.
-Rev.
-Applied 17, 064017 (2022).
-- [36] Zhong, J. et al. Observation of Anomalous Planar Hall Effect Induced by One-Dimensional Weak Antilocalization. ACS Nano acsnano.3c10120 (2024)
-- [37] L. Qiao et al., Ultralong Single-Crystal -Bi4Br4 Nanobelts with a High Current Carrying Capacity by Mechanical Exfoliation, J. Phys. Chem. C 125, 22312 (2021).
-- [38] S.-L.
-Wu, Z.-H.
-Ren, Y.-Q.
-Zhang, Y.-K.
-Li, J.-F.
-Han, J.-X.
-Duan, Z.-W.
-Wang, C.-Z.
-Li, and Y.-G.
-Yao, Gate-tunable transport in van der Waals topological insulator Bi4Br4 nanobelts, J.
-Phys.: Condens.
-Matter 35, 234001 (2023).
-- [39] M. S. Hossain et al., Quantum transport response of topological hinge modes, Nat. Phys. 1 (2024).
-- [40] R. Noguchi et al., A weak topological insulator state in quasi-one-dimensional bismuth iodide, Nature 566, 7745 (2019).
-- [41] Bapat, R. B, Resistance matrix of a weighted graph, Communications in Mathematical and in Computer Chemistry 50, 73‑82 (2004).
-- [42] Devriendt Karel, Graph Geometry from Effective Resistances, PhD Thesis, Trinity College, 2022.
-- [43] Datta, S. Electronic Transport in Mesoscopic Systems. (Cambridge Univ. Press, Cambridge, 2009).
-- [44] E. Akkermans and G. Montambaux, Mesoscopic Physics of Electrons and Photons, Cambridge University Press (2007).
-- [45] S. Hikami, A. I. Larkin, and Y. Nagaoka, Spin-Orbit Interaction and Magnetoresistance in the Two Dimensional Random System, Progress of Theoretical Physics 63, 707 (1980).
-- [46] H.
-Steinberg, J.-B.
-Laloë, V.
-Fatemi, J.
-S.
-Moodera, and P.
-Jarillo-Herrero, Electrically tunable surface-to-bulk coherent coupling in topological insulator thin films, Phys.
-Rev.
-B 84, 233101 (2011).
-- [47] J.
-Chen, X.
-Y.
-He, K.
-H.
-Wu, Z.
-Q.
-Ji, L.
-Lu, J.
-R.
-Shi, J.
-H.
-Smet, and Y.
-Q.
-Li, Tunable surface conductivity in Bi 2 Se 3 revealed in diffusive electron transport, Phys.
-Rev.
-B 83, 241304 (2011).
-- [48] P. A. Lee, A. D. Stone, and H. Fukuyama, Universal conductance fluctuations in metals: Effects of finite temperature, interactions, and magnetic field, Phys. Rev. B 35, 1039 (1987).
-- [49] P. Mohanty and R. A. Webb, High-Field Measurements of Electron Decoherence Time in Metallic Nanowires: Switching off Magnetic Impurity Spins, Phys. Rev. Lett. 91, 066604 (2003).
-- [50] C. W. J. Beenakker and H. Van Houten, Flux-cancellation effect on narrow-channel magnetoresistance fluctuations, Phys. Rev. B 37, 6544 (1988).
-- [51] Groth, C. W., Wimmer, M., Akhmerov, A. R. and Waintal, X. Kwant: a software package for quantum transport. New J. Phys. 16, 063065 (2014).
-- [52] M. Büttiker, Absence of backscattering in the quantum Hall effect in multiprobe conductors, Phys. Rev. B 38, 9375 (1988).
-- [53] M. Büttiker, Symmetry of electrical conduction, IBM J. Res. Dev. 32, 317 (1988).
-- [54] B. Gao, Y. F. Chen, M. S. Fuhrer, D. C. Glattli, and A. Bachtold, Four-Point Resistance of Individual Single-Wall Carbon Nanotubes, Phys. Rev. Lett. 95, 196802 (2005).
-- [55] G. Timp, A. M. Chang, P. Mankiewich, R. Behringer, J. E. Cunningham, T. Y. Chang, and R. E. Howard, Quantum transport in an electron-wave guide, Phys. Rev. Lett. 59, 732 (1987).
-- [56] de Picciotto, R., Stormer, H. L., Pfeiffer, L. N., Baldwin, K. W. and West, K. W. Four-terminal resistance of a ballistic quantum wire. Nature 411, 51–54 (2001).
-- [57] S. Washburn and R. A. Webb, Aharonov-Bohm effect in normal metal quantum coherence and transport, Advances in Physics 35, 375 (1986).
-- [58] Altshuler, B. L., Aronov, A. G. and Spivak, B. Z. The Aharonov–Bohm effect in disordered conductors. JETP Lett. 33, 94–97 (1981).
-- [59] Washburn, S., Umbach, C. P., Laibowitz, R. B. and Webb, R. A. Temperature dependence of the normal-metal Aharonov–Bohm effect. Phys. Rev. B 32, 4789–4792 (1985).
-- [60] Dufouleur, J. et al. Quasiballistic Transport of Dirac Fermions in a Bi 2 Se 3 Nanowire. Phys. Rev. Lett. 110, 186806 (2013).
-- [61] Jauregui, L.
-A., Pettes, M.
-T., Rokhinson, L.
-P., Shi, L.
-and Chen, Y.
-P.
-Magnetic field-induced helical mode and topological transitions in a topological insulator nanoribbon.
-Nature Nanotech 11, 345–351 (2016).
-- [62] G. Behner et al., Aharonov-Bohm Interference and Phase-Coherent Surface-State Transport in Topological Insulator Rings, Nano Lett. 23, 6347 (2023).
-- [63] Pan, Z.-C. et al. Altshuler-Aronov-Spivak interference of one-dimensional helical edge states in . Phys. Rev. B 107, 045411 (2023).
-- [64] A. E. Hansen, A. Kristensen, S. Pedersen, C. B. Sørensen, and P. E. Lindelof, Mesoscopic decoherence in Aharonov-Bohm rings, Phys. Rev. B 64, 045327 (2001).
-- [65] See Supplemental Material for a discussion of the crystalline quality of Bi4Br4 flakes, resistance networks of two other samples at different temperatures, the length dependence of resistance for samples and , we expand on the temperature dependence of the Aharonov-Bohm effect for sample , details of the numerical simulations presented in the main text, the fitting of the magnetoconductance using the formula for 1D weak anti-localization, and finally a presentation of zero-bias anomaly data and comparison with theoretical predictions.
+This work was funded by JSPS KAKENHI (Grants No.
+21H04652, No.
+21K18181 and No.
+21H05236), JST CREST (Grant No.
+JPMJCR20B4), the French Renatech network, the CPER Hauts de France project IMITECH, the Métropole Européenne de Lille, and the European Union through the BALLISTOP ERC 66566 advanced grant.DATA AVAILABILITY
+The data that support the findings of this article are openly available [86].